News & Status
  • News & status not loaded yet.

Javascript is turned off in your web browser.

To take full advantage of this site, please turn it on and then refresh the page.

Microscope Facility News & Equipment Status

To change News view:
Two months
Quarter
Year
All
Select multiple with control-key
Restrict listing
to selected
equipment
(Exclude announcements)

2017-09-19 (Zeiss LSM 510-meta #2): Zeiss LSM 510-Meta (#2) >> Back online, safety bypass switch was the issue

9/20/2017 4:01pm (UPDATE)>> Back online, safety bypass switch was the issue
The 510 system has been serviced and is back online. Thank you for your patience.

****************************************
9/19/2017 2:12pm Zeiss LSM 510 Meta #2 Not Available
The Zeiss LSM 510 Meta #2 is not available until further notice due to a broken main shutter component. We will notify all users once we have an ETA on repairs.

2017-09-18 (Zeiss LSM 780 FCS): UPDATE >> Replacement argon laser for 780 has arrived, installation scheduled for Mon Sept 18

9/18/2017 9:41am (UPDATE)>>The replacement argon laser for the 780 will be installed the Zeiss engineer today. Thank you for your patience.

****************************************
8/30/2017 9:12am (UPDATE)>>As of today, the argon laser on the 780 has become intermittent and therefore has been turned off. Any samples that need to be excited with 457, 477, 488 and 514 (i.e. CFP, GFP, YFP) will not work on the 780.

The new laser is en route but delayed in customs. We apologize for the inconvenience and will send another update as soon as we have more information.

****************************************
8/21/2017 5:00pm>>The replacement part is backordered, however, the Zeiss 780 remains online and available for use. Thank you for your patience!

****************************************
8/16/2017 2:12pm (UPDATE)>>780 remains online and available for use
The 780 will remain online and available for use, but with a couple of caveats. Here are some comments/guidelines from Hoku:
- All other laser lines are perfectly functional and require no changes to your current protocols. Anything using the 405, 561, and 633nm lines should be completely unperturbed.
- The 514 line is too degraded to use (unstable power, low output). This means that CFP/YFP FRET experiments are not viable until the new laser head arrives. Furthermore, if you are imaging YFP alone, use the 488 line. We can show you how to modify your protocols.
- The 457 and 488 lines are still working, but at reduced power. You may need to increase the laser power on your specimens to obtain an adequate image in green or CFP.
- If you are doing quantitative imaging using the argon-ion laser (such as measuring raw intensities across specimens), you may wish to wait until the new laser head arrives. However, for most uses (such as determining the localization of a protein, monitoring cell movement or morphology, etc), you will be able to continue to use the system uninterrupted, turning the power up for now and turning it back down again when the new laser arrives.
If you have any questions, please don’t hesitate to ask one of our staff members. Thank you for your patience!

****************************************
8/16/2017 2:11pm>>Argon laser is off (CFP/GFP/YFP will not work), remains online, awaiting replacement laser.

The argon-ion laser on the LSM 780 (457, 488, and 514nm lines) is failing. A replacement has been ordered, but it will be a few days before it gets in. We apologize for the inconvenience.

2017-09-15 (Philips CM120 TEM): New Rules (repaired specimen holder)

In response to abuse/breaking the Philips specimen holder:
Effective Mon Sept 18, 2017:

  1. Do NOT use the specimen holder without re-training
  2. EITHER:
    SCHEDULE a short re-training (<15min, FREE),   OR
    WAIT for staff availability to re-train you at the start of your session.
  3. We will SUSPEND/RESTRICT future privileges of anyone operating the instrument without re-training after 9/18/17.

2017-09-05 (Philips CM120 TEM): Philips CM120 TEM Not Available: Grid Holder Broken

The Philips CM120 TEM is not available for use until further notice. The grid holder is broken. Please be careful with the microscope! Carelessness is COSTLY! We now have to wait for this one to be repaired or a replacement ordered. We will email and post a news item as soon as we have new information.

2017-08-29 (3i Spinning Disk Confocal,Leica Cryostat,Leica EM ICE High Pressure Freezer): Physiol G05 Plumbing Repair

9-5-17 11:00AM As of today, all equipment in G05 is functioning and available.

The sewage/drain lines in the ceiling of Physiology G05 have to be replaced. In preparation, we have to shut down ALL equipment in the room: HPF, cryostat, ultramicrotome, chromium evaporator, 3i spinning disk and rotary evaporator. Expected down time for equipment is Tue 8/29 to Tue 9/5.

Thank you for your understanding.

2017-07-25 (Zeiss LSM 780 FCS): Zeiss LSM780 BACK ONLINE: part replaced, system recalibrated

7/25/2017 9:28am (UPDATE)>>
As of 9:30am, the 780 is back online. All user reservations from Wed, July 19 to Monday, July 24 have been canceled. Thank you for your patience!

****************************************

7/24/2017 4:34pm (UPDATE)>>Zeiss LSM780 now in repair, down until Tuesday 7/25
The Zeiss engineer has replaced the part on the 780, but as of 4:30pm, he is still recalibrating the instrument. Even if he is finished by tonight, Barbara or Hoku will have to inspect the system tomorrow morning before we can bring the 780 back online for our users.

****************************************

7/20/2017 11:56am (UPDATE)>>
The 780 will be down until Monday. Unfortunately, that is the earliest date that we can expect to receive parts and service. We really appreciate your patience.

****************************************

7/19/2017 2:42pm
The Zeiss LSM 780 is down. Multiple errors are appearing in the software but it may be a hardware issue. We are awaiting a phone call/visit from Zeiss and we will try to keep you updated. We will cancel all reservations from today so please do not be concerned that you will be charged for unused time. Thank you for your patience.

2017-07-05: New server TempStorage4

The Microscope Facility has a new server for its users: TempStorage4 (>3x bigger & >6x faster). Please delete your shortcut to the old server and create a shortcut to the new one at \\162.129.230.229\TempStorage4 for Windows and //162.129.230.229/TempStorage4 for Macs (see instructions if you've forgotten how). The login and password to the new server is the same as before. Please copy your data from the old server (accessible as a read-only subdirectory 'TempStorage3' on TempStorage4) to the new server.

6/30/2017 11:41am>>
TempStorage3 Offline/to Retire
TempStorage3 is offline while we bring up the new Server (Tempstorage4), which is >3x bigger and >6x faster. You will need to edit your shortcuts to the new folder (Windows: "\\162.129.230.229\Tempstorage4"; Mac: "162.129.230.229/Tempstorage4"), noting that the 's' is lowercase. The login/password is the same as before for TempStorage3, but will no longer be accessible.

If you don't know how to create such shortcuts, see FastFacts/Network.

Direct access to TempStorage3 (SMB1 or FTP) will no longer be available. Your old data will still be accessible through the new server as a linked folder called 'TempStorage3' and you can copy your old data folder from the old server to the new server. Eventually, TempStorage3 will shut down and retire (linked folder will disappear), but you'll have at least one month to transfer your data.


Last modified